OpenStudy (anonymous):

Using the quadratic formula identify the a, b and c values from the equation: x^2+4x+1=0

OpenStudy (calculusxy):

The quadratic formula is (x+a)(x+b)=x^2+bx+ax+ab=x^2+(ab)^2+ab

OpenStudy (anonymous):

I got a:1 B:4 C:1

OpenStudy (calculusxy):

You have to reorder the terms.

OpenStudy (calculusxy):

It will make it: 1+4x+x^2=0

OpenStudy (calculusxy):

Did you find your answer?